Question 3 10 Points A garage door control system has the following inputs: A- A physical button (A), which outputs 1 whiAbro B
Answer #1

Below screenshot gives you the required solution.

Q3.) As per table will be given condition, truth A B & Q= A +6 0 1 1 1 01 1 1 Now, we will check all the given give above truoption A = (A+B) + A+B) A+B Option B A D -Q = (A+B) * 8 = AB+B - ABS (A+1) B 2 В e to Am option D -Q = Ate D- Q = 476 DA QE A

So out of all the given options, option A, C and F give the correct design for the circuit.

Hope it helps, give it a thumbs up.

Feel free to comment in case of any query.